# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 5. Discrete two-colored dual hive model
Without loss of generality, suppose that $e\in E_{n}$ has type $\ell$ and let $s$ be a lozenge such that $e$ is a border edge of type $\ell$ of $s$ and the opposite edge $e^{\prime}$ is a translation of $e$ such that $h(e)\geq h(e^{\prime})$ and $e_{\ell+1}\leq e^{\prime}_{\ell+1}$. Hence, either $h(e)=h(e^{\prime})$, $e_{\ell+1}=e^{\prime}_{\ell+1}-1$ and the middle edge $f$ of $s$ is of type $k=\ell-1$ or $h(e^{\prime})=h(e)-1$, $e^{\prime}_{\ell+1}=e_{\ell+1}$ and the middle edge $f$ is of type $k=\ell+1$. Suppose now that at least one of the edges of the lozenge is colored $m$ and the middle edge $f$ is not dual to an edge of $\hat{\mathcal{G}}_{P}$ coming from the blowing-up of an even vertex. Hence, $f$ corresponds to a strip $S_{f}$ of $k$. Moreover, there exist $\tilde{e}$ and $\tilde{e^{\prime}}$ of type $\ell$ with $h(\tilde{e})=L(e)$ and $h(\tilde{e}^{\prime})=L(e^{\prime})$, $\tilde{f},\tilde{f^{\prime}}\in S_{f}$ of type $k$ such that $\tilde{e}$ and $\tilde{f}$ (resp. $\tilde{e}^{\prime}$ and $\tilde{f}^{\prime}$) comes from a same vertex $v$ of $P_{v}$ (either directly or after a blowing-up). Then $S_{f}$ corresponds to a genuine strip $S_{f}$ with ending edges $\tilde{f},\tilde{f^{\prime}}$ and there exists $\tilde{e},\tilde{e}^{\prime}$ of type $\ell$ with $h(\tilde{e})=L(e)$, $h(\tilde{e}^{\prime})=L(e^{\prime})$ and $\tilde{e}$ (resp. $\tilde{e}^{\prime}$) belongs to a same vertex as $\tilde{f}$ (resp. $\tilde{f}^{\prime}$) in $P_{v}$. Remark that $\tilde{f}_{\ell}\geq\tilde{f}^{\prime}_{\ell}$ for otherwise there would an edge of type different from $f$ and labeled 00 or $1$ in the strip $S_{f}$, which is not possible from Figure 10. Then, if $\tilde{e}$ is of type $\ell$ and $\tilde{f}$ is of type $\ell-1$ coming from a same triangle of $T_{n}$, we have the following relation between $\tilde{f}_{\ell}$ and $h(\tilde{e})$ depending on the orientation of the triangle and the colors of the boundary edges (the color and position of $\tilde{f}$ is bold)
@@
Figure 11. Coordinates of an edge in fonction of the coloring and height of the next edge in a triangle ($\tilde{f}$ correspond to the bold edge and $\tilde{e}$ corresponds to the horizontal edge).
From those relation and the fact that $\tilde{f}_{\ell}\geq\tilde{f}^{\prime}_{\ell}$, we deduce that $h(\tilde{e})\geq h(\tilde{e}^{\prime})$ in the case $e^{\prime}_{\ell+1}=e_{\ell+1}+1$ and that $h(\tilde{e})>h(\tilde{e}^{\prime})$ if one of the boundary edge of $s$ different from $e$ is colored $m$. The case $e_{\ell}=e^{\prime}_{\ell}+1$ is done similarly, yielding always $h(\tilde{e})>h(\tilde{e}^{\prime})$.
Finally, if the middle edge is coming from the blowing up of an even vertex and is not colored $m$, then this edge is necessarily colored $1$, and thus $e$ and $e^{\prime}$ are colored 00 and the opposite edges are colored $1$. The strict inequality is directly deduced from construction of $m$ and Lemma 4.7 giving the height of edges colored 00 in an even vertex. ∎
We deduce that $T_{n}$ with the labelling $(C,L)$ is a genuine discrete two-colored dual hive, which we denote by $\Phi(P)$.
Proposition 5.8.
The map $\Phi$ is bijective.
Proof.
Let us construct the reverse bijection. Let $H$ be a two-colored dual hive. We first define the candidate vertex partition (without the coloring for now) $\mathcal{P}_{v}$ as follows :
  • for each triangle face $t=(t^{0},t^{1},t^{2})\in T_{n}$, with $t^{\ell}$ of type $\ell$, with boundary colors in $(0,0,0)$, $(1,1,1)$ or $(1,0,3)$ (up to a cyclic order), we define a block $B_{t}\subset E_{N}$ with edges $e^{0},e^{1},e^{0}$, with $e^{\ell}$ of type $\ell$, and such that :
@@
  • for each edge $s$ of type $\ell$ on the boundary of $T_{n}$, we define a singleton in $B_{e}\subset E_{N}$ consisting of the unique edge $e$ of type $\ell$ with $h(e)=L(s)$.
Moreover, by Lemma 5.9 below, edges $e,e^{\prime}$ coming from different edge $t,t^{\prime}\in T_{n}$ by the previous constructions are distinct. Let $\mathcal{E}=\bigcup_{B\in\mathcal{P}_{v}}B$ We can thus define a coloring $c$ on $\mathcal{E}$ by setting $c(e)=C(t)$ when $e$ is constructed from $t$ above. By construction and the property $(1)$ of Definition 5.2, the covering $\mathcal{P}_{v}$ satisfies all the properties of Lemma 4.7. With the above constructions and the conditions (1) and (2.a) of a two-color dual hive, we can then check that all the relations from Figure 11 is still satisfied when the bold edge is an element of $\mathcal{E}$.
Define then a relation $\sim$ on $\mathcal{E}$ by saying that $e\sim e^{\prime}$ if $e,e^{\prime}$ are coming from a same edge of $T_{n}$ through the previous construction, and denote by $\mathcal{P}_{e}$ the set partition coming from this relation. By the properties of a two-colored dual hive, any long rhombus considered before of border edges of type $\ell,\ell+1$ has its inner middle edges of type $\ell-1$ colored $m$, so that none of the triangles inside this long rhombus yields block of $\mathcal{P}_{v}$ through the first step. Hence, each edge of $T_{n}$ yields at most $2$ edges of $\mathcal{E}$. Remark that an edge of $T_{n}$ is either adjacent to two faces or to one face and the boundary of $T_{n}$, so that $\mathcal{P}_{e}$ consists of pairs or singleton, and in the latter case the singleton belongs to two blocks of $\mathcal{P}_{v}$. If $e\sim e^{\prime}$, by the above construction $c(e)=c(e^{\prime}),L(e)=L(e^{\prime})$ and $e,e^{\prime}$ have a same type, so that $\mathcal{P}_{e}$ only has admissible pairs (which can be reduced to a singleton). If $e\not=e^{\prime}$ belong to a same block of $\mathcal{P}_{v}$ they come from different edges of $T_{n}$ and thus $\{e,e^{\prime}\}\not\in\mathcal{P}_{e}$.
@@
Pairs of $\mathcal{P}_{e}$ are admissible and any different pairs $B,B^{\prime}\in\mathcal{P}_{e}$ do not cross, thus partition $\mathcal{P}_{e}$ satisfies the properties of Lemma 4.9. Finally, by Proposition 4.11 applied to $(\mathcal{P}_{v},\mathcal{P}_{e})$, there exists a unique puzzle $P$ such that the corresponding vertex and edge partitions are respectively $\mathcal{P}_{v}$ and $\mathcal{P}_{e}$. Denote by $\Psi(H)$ this puzzle. It is then straightforward to prove that $\Psi\circ\Phi$ and $\Phi\circ\Psi$ are identity maps. ∎
Lemma 5.9.
Let $t\not=t^{\prime}\in T_{n}$ of same type $\ell$ in blocks of $\mathcal{P}_{v}$, and suppose without loss of generality that $t_{\ell+1}>t^{\prime}_{\ell+1}$ or $t_{\ell+1}=t^{\prime}_{\ell+1}$ and $h(t^{\prime})>h(t)$. Then, if $h(t)<h(t^{\prime})$ and $t_{\ell+1}>t^{\prime}_{\ell+1}$, the edges $e^{1},e^{2}$ (resp. $e^{3},e^{4}$) of $T_{N}$ associated to $t$ (resp. $t^{\prime}$) satisfy
@@
If $t_{\ell+1}=t^{\prime}_{\ell+1}$ and $h(t^{\prime})>h(t)$, then by Condition $(2.c)$ of Definition 5.2, $L(t^{\prime})>L(t)$ except if the middle edge of all lozenges between $t$ and $t^{\prime}$ are colored $m$. In the latter case, let $s$ (resp. $s^{\prime}$) be the edge of type $\ell-1$ such that $t,s$ form a reverse triangle (resp. $t^{\prime},s^{\prime}$ form a direct triangle). Since $s_{\ell}=t_{\ell}+1$ and $s^{\prime}_{\ell}=t^{\prime}_{\ell}$, the inequality $t^{\prime}_{\ell}>t_{\ell}$ implies that $s^{\prime}_{\ell}>s_{\ell}$. Similarly, since $s_{\ell-1}=t_{\ell-1}-1$ and $s^{\prime}_{\ell-1}=t^{\prime}_{\ell-1}$, we have
@@
Hence, $L(s^{\prime})\leq L(s)$. Let us introduce the third edge $r$ (resp. $r^{\prime}$) of the triangle with edges $s,t$ (resp. $s^{\prime},t^{\prime})$. By Figure 11 and Condition (1) from Definition 5.2, we get that $e^{3}_{\ell-1}=N-e^{3}_{\ell}-e^{3}_{\ell+1}=N-L(t^{\prime})-L(r^{\prime})-1=L(s^{\prime})$ and $e^{2}_{\ell-1}=N-e^{2}_{\ell}-e^{2}_{\ell+1}=N-L(t)-L(r)=L(s)+1$ and thus $e^{2}_{\ell-1}=L(s)+1>L(s)=e^{3}_{\ell-1}$, so that
@@
If $h(t)\geq h(t^{\prime})$ and $t_{\ell+1}>t^{\prime}_{\ell+1}$, then $t^{\prime}_{\ell+2}>t_{\ell+2}$. Suppose without loss of generality that $e^{1}_{\ell+1}\geq e^{2}_{\ell+1}$ and $e^{3}_{\ell+1}\geq e^{4}_{\ell+1}$. Let us consider the edges $s,s^{\prime}$ of type $\ell+1$ such that $(t,s,r)$ and $(t^{\prime},s^{\prime},r^{\prime})$ are respectively direct and reverse triangles of $T_{n}$ so that the corresponding edge of $t$ and the piece containing the direct triangle is $e^{2}$ and the corresponding edge for $t^{\prime}$ and the reverse triangle is $e^{3}$. Since $h(s)=t_{\ell+1}-1$ and $s_{\ell+2}=t_{\ell+2}+1$ and $h(s^{\prime})=t^{\prime}_{\ell+1}-1$ and $s^{\prime}_{\ell+2}=t^{\prime}_{\ell+2}$, $h(s)>h(s^{\prime})$ and $s^{\prime}_{\ell+2}\geq s_{\ell+2}$, so that $L(s)>L(s^{\prime})$ by Remark 5.3. Then, since $(t,u,s)$ is a direct triangle, Figure 11, $e^{2}_{\ell+1}=L(s)+1$, except if $c(t)=1,c(s)=0$ and $c(u)=m$ where $e^{2}_{\ell+1}=L(s)+2$. Likewise, since $(t^{\prime},s^{\prime},u^{\prime})$ is a reverse triangle, $e^{3}_{\ell+1}=L(s^{\prime})+1$ except if $c(t)=1,c(s)=0,c(u)=m$ or $c(t)=0,c(s)=m,c(u)=1$ where $e^{3}_{\ell+1}=L(s^{\prime})$. Hence, in any case,
@@
Lemma 5.10.
Let $t,t^{\prime}\in T_{n}$ be of respective type $\ell,\ell+1$ yielding edges in $\mathcal{E}$, and denote by $e^{1},e^{2}$ (resp. $e^{3},e^{4}$) the edges of $T_{N}$ corresponding to $t$ (resp. $t^{\prime}$). Then,
@@
The proof of the four assertion are similar.
@@
    Suppose that $t^{\prime}_{\ell}>t_{\ell}$ and $t^{\prime}_{\ell+1}<t_{\ell+1}$. Let $(t^{\prime},s^{\prime},r^{\prime})$ and $(t^{\prime},s^{\prime\prime},r^{\prime\prime})$ be the reverse and direct triangle belonging to pieces yielding respectively $e^{3}$ and $e^{4}$, with $s^{\prime},s^{\prime\prime}$ of type $\ell$. Since then $e^{4}_{\ell}\geq e^{3}_{\ell}$, it suffices to show that $e^{3}_{\ell}>e^{2}_{\ell}$. Since $s^{\prime}_{\ell}=t^{\prime}_{\ell}-1$ and $s^{\prime}_{\ell+1}=t^{\prime}_{\ell+1}+1$, we have $s^{\prime}_{\ell}\geq t_{\ell}$ and $t_{\ell+1}\geq s^{\prime}_{\ell+1}$, so that $L(s^{\prime})\geq L(t)$. Since $t^{\prime}$ is of type $\ell+1$ not and colored $m$ and $(t^{\prime},s^{\prime},r^{\prime})$ is a reverse triangle, Figure 11 and Condition (1) from Definition 5.2 yield that either $e^{3}_{\ell+2}=L(r^{\prime})+1$ and $L(r^{\prime})+L(s^{\prime})+L(t^{\prime})=N-2$ or $e^{3}_{\ell+2}=L(r^{\prime})$ and $L(r^{\prime})+L(s^{\prime})+L(t^{\prime})=N-1$. In any case, $e^{3}_{\ell+2}=N-1-L(t^{\prime})-L(s^{\prime})$, so that
    $e^{3}_{\ell}=N-e^{3}_{\ell+1}-e^{3}_{\ell+2}=N-L(t^{\prime})-(N-1-L(r^{\prime})-L(s^{\prime})=L(s^{\prime})+1>L(t)=e^{2}_{\ell}=e^{1}_{\ell}.$
@@
    Suppose that $t^{\prime}_{\ell}<t_{\ell}$ and $t^{\prime}_{\ell+1}\geq t_{\ell+1}$, and let $s^{\prime}$ be the edge of type $\ell$ such that $(t^{\prime},s^{\prime},r^{\prime})$ is a direct triangle. Since $s^{\prime}_{\ell}=t^{\prime}_{\ell}$ and $s^{\prime}_{\ell+1}=t^{\prime}_{\ell+1}+1$, $s^{\prime}_{\ell}<t_{\ell}$ and $s^{\prime}_{\ell+1}>t_{\ell+1}$, so that $L(t)>L(s^{\prime})$ by Remark 5.3. Since $(t^{\prime},s^{\prime},r^{\prime})$ is a direct triange, Figure 11 and Condition (1) from Definition 5.2 yield by a same reasoning as above that $e^{4}_{\ell+2}\geq L(r^{\prime})+1=N-L(s^{\prime})-L(t^{\prime})$, so that, using that $e^{4}_{\ell+1}=L(t^{\prime})$,
@@
    Suppose that $t^{\prime}_{\ell}\geq t_{\ell}$ and $t^{\prime}_{\ell+1}\geq t_{\ell+1}$, and let $s$ be the edge of type $\ell+1$ such that $(t,s,r)$ is a reverse triangle. Then, $s_{\ell}=t_{\ell}+1$ and $s_{\ell+1}=t_{\ell+1}-1$. Hence, $t^{\prime}_{\ell+1}>s_{\ell+1}$ and $t^{\prime}_{\ell+2}=N-t^{\prime}_{\ell+1}-t^{\prime}_{\ell}\leq N-s_{\ell+1}-1-s_{\ell}+1\leq s_{\ell+2}$ and the inequality is strict except when $t^{\prime}_{\ell}=t_{\ell}$ and $t^{\prime}_{\ell+1}=s_{\ell+1}$. Hence, by Remark 5.3 ins the case of strict inequality and Condition (2.b) and (2.c) from Definition 5.2, $L(t^{\prime})>L(s)$, except when $t^{\prime}_{\ell}=t_{\ell},t^{\prime}_{\ell+1}=s_{\ell+1}$ and $C(u)=m$, in which case $L(t^{\prime})=L(s)$. In the first case, by Figure 11 we have $e^{2}_{\ell+1}\leq L(s)+1\leq L(t^{\prime})$. In the second case, since $C(r)=m$ we have $e^{2}_{\ell+1}=L(s)\leq L(t^{\prime})$, so that in any case

### v2

@@ 5. Discrete two-colored dual hive model
Without loss of generality, suppose that $e\in E_{k}$ has type $\ell$ and let $s$ be a lozenge such that $e$ is a border edge of type $\ell$ of $s$ and the opposite edge $e^{\prime}$ is a translation of $e$ such that $h(e)\geq h(e^{\prime})$ and $e_{\ell+1}\leq e^{\prime}_{\ell+1}$. Hence, either $h(e)=h(e^{\prime})$, $e_{\ell+1}=e^{\prime}_{\ell+1}-1$ and the middle edge $f$ of $s$ is of type $k=\ell-1$ or $h(e^{\prime})=h(e)-1$, $e^{\prime}_{\ell+1}=e_{\ell+1}$ and the middle edge $f$ is of type $k=\ell+1$. Suppose now that at least one of the edges of the lozenge is colored $m$ and the middle edge $f$ is not dual to an edge of $\hat{\mathcal{G}}_{P}$ coming from blowing-up an even vertex. Hence, $f$ corresponds to a strip $S_{f}$ of type $k$. Moreover, there exist $\tilde{e},\tilde{e^{\prime}}\in E_{N}^{2}$ of type $\ell$ with $h(\tilde{e})=L(e)$ and $h(\tilde{e}^{\prime})=L(e^{\prime})$, $\tilde{f},\tilde{f^{\prime}}\in S_{f}$ of type $k$ such that $\tilde{e}$ and $\tilde{f}$ (resp. $\tilde{e}^{\prime}$ and $\tilde{f}^{\prime}$) comes from a same vertex $v$ (resp. $v^{\prime}$) of $P_{v}$ (either directly or after a blowing-up).
Remark that $\tilde{f}_{\ell}\geq\tilde{f}^{\prime}_{\ell}$ for otherwise, in the strip $S_{f}$, there would be an edge of type different from $f$ and labeled 00 or $1$ coming from $v$, which is not possible from Figure 10. Then, if $\tilde{e}$ is of type $\ell$ and $\tilde{f}$ is of type $\ell-1$ coming from a same triangle of $T_{N}$, we resume in Figure 12 the relation between $\tilde{f}_{\ell}$ and $h(\tilde{e})$ depending on the orientation of the triangle and the colors of the boundary edges (the color and position of $\tilde{f}$ is bold). Those relations are consequences of Lemma 4.9 and blowups of even degree vertices.
@@
Figure 12. Coordinates of an edge in function of the coloring and height of the next edge in a triangle ($\tilde{f}$ correspond to the bold edge and $\tilde{e}$ corresponds to the horizontal edge).
From those relation and the fact that $\tilde{f}_{\ell}\geq\tilde{f}^{\prime}_{\ell}$, we deduce that $h(\tilde{e})\geq h(\tilde{e}^{\prime})$, i.e $L(e)\geq L(e^{\prime})$, in the case $e^{\prime}_{\ell+1}=e_{\ell+1}+1$ and that $h(\tilde{e})>h(\tilde{e}^{\prime})$, i.e $L(e)>L(e^{\prime})$, if one of the boundary edge of $s$ different from $e$ is colored $m$. The case $e_{\ell}=e^{\prime}_{\ell}+1$ is done similarly, yielding always $h(\tilde{e})>h(\tilde{e}^{\prime})$, i.e $L(e)>L(e^{\prime})$.
Finally, if the middle edge is coming from the blowing up of an even vertex and is not colored $m$, then this edge is necessarily colored $1$, and thus $e$ and $e^{\prime}$ are colored 00 and the opposite edges of their lozenge are colored $1$. The strict inequality is directly deduced from construction of $L=h\circ\zeta^{-1}$ and Lemma 4.9 giving the height of edges colored 00 in an even vertex.
Hence, $T_{k}$ with the labelling $(C,L)$ is a genuine discrete two-colored dual hive, which we denote by $\zeta(P)$. The boundary values are directly deduced from the boundary of $P$, so that the resulting hive is in $\mathcal{H}(c(\omega),l(\omega),N)$. ∎
Proof of Theorem 5.3.
Let us construct the reverse bijection. Let $H=(C,L)$ be a two-colored dual hive in $\mathcal{H}(c(\omega),l(\omega),N)$. We first define the candidate vertex partition (without the coloring for now) $\mathcal{P}_{v}$ as follows :
  • for each triangle face $t=(t^{0},t^{1},t^{2})\in T_{k}$, with $t^{\ell}$ of type $\ell$, with boundary colors in $(0,0,0)$, $(1,1,1)$ or $(1,0,3)$ (up to a cyclic order), we define a block $B_{t}\subset E_{N}$ with edges $e^{0},e^{1},e^{0}$, with $e^{\ell}$ of type $\ell$, and such that :
@@
  • for each edge $s$ of type $\ell$ on the boundary of $T_{k}$, we define a singleton in $B_{e}\subset E_{N}$ consisting of the unique edge $e$ of type $\ell$ with $h(e)=L(s)$.
Moreover, by Lemma 5.10 below, edges $e,e^{\prime}$ coming from different edges $t,t^{\prime}\in T_{k}$ by the previous constructions are distinct. Let $\mathcal{E}=\bigcup_{B\in\mathcal{P}_{v}}B$. We can thus define a coloring $c$ on $\mathcal{E}$ by setting $c(e)=C(t)$ when $e$ is constructed from $t$ above. By construction and the property $(1)$ of Definition 5.1, the covering $\mathcal{P}_{v}$ satisfies all the properties of Lemma 4.9. With the above constructions and the conditions (1) and (2.a) of a two-color dual hive, we can then check that all the relations from Figure 12 is still satisfied when the bold edge is an element of $\mathcal{E}$.
Define then a relation $\sim$ on $\mathcal{E}$ by saying that $e\sim e^{\prime}$ if $e,e^{\prime}$ are coming from a same edge of $T_{k}$ through the previous construction, and denote by $\mathcal{P}_{e}$ the set partition coming from this relation. By the properties of a two-colored dual hive, any long rhombus considered before of border edges of type $\ell,\ell+1$ has its inner middle edges of type $\ell-1$ colored $m$, so that none of the triangles inside this long rhombus yields block of $\mathcal{P}_{v}$ through the first step. Hence, each edge of $T_{k}$ yields at most $2$ edges of $\mathcal{E}$. Remark that an edge of $T_{k}$ is either adjacent to two faces or to one face and the boundary of $T_{k}$, so that $\mathcal{P}_{e}$ consists of pairs or singletons, and in the latter case the singleton belongs to two blocks of $\mathcal{P}_{v}$. If $e\sim e^{\prime}$, by the above construction $c(e)=c(e^{\prime}),L(e)=L(e^{\prime})$ and $e,e^{\prime}$ have the same type, so that $\mathcal{P}_{e}$ only has admissible pairs (which can be reduced to a singleton). If $e\not=e^{\prime}$ belong to a same block of $\mathcal{P}_{v}$ they come from different edges of $T_{k}$ and thus $\{e,e^{\prime}\}\not\in\mathcal{P}_{e}$.
@@
Pairs of $\mathcal{P}_{e}$ are admissible and any two different pairs $B,B^{\prime}\in\mathcal{P}_{e}$ do not cross, thus partition $\mathcal{P}_{e}$ satisfies the properties of Lemma 4.11. Finally, by Proposition 4.14 applied to $(\mathcal{P}_{v},\mathcal{P}_{e})$, there exists a unique puzzle $P$ such that the corresponding vertex and edge partitions are respectively $\mathcal{P}_{v}$ and $\mathcal{P}_{e}$. Denote by $\chi(H)$ this puzzle. It is clear from the above constructions that $\chi(H)\in P(\omega)$ and that $\chi\circ\zeta$ and $\zeta\circ\chi$ are respectively identity maps of $P(\omega)$ and $\mathcal{H}(c(\omega),l(\omega),N)$. ∎
Lemma 5.10 (Same type blocks do not cross).
Let $t\not=t^{\prime}\in T_{k}$ of same type $\ell$ in blocks of $\mathcal{P}_{v}$, and suppose without loss of generality that $t_{\ell+1}>t^{\prime}_{\ell+1}$ or $t_{\ell+1}=t^{\prime}_{\ell+1}$ and $h(t^{\prime})>h(t)$. Then, if $h(t)<h(t^{\prime})$ and $t_{\ell+1}>t^{\prime}_{\ell+1}$, the edges $e^{1},e^{2}$ (resp. $e^{3},e^{4}$) of $T_{N}$ associated to $t$ (resp. $t^{\prime}$) satisfy
@@
If $t_{\ell+1}=t^{\prime}_{\ell+1}$ and $h(t^{\prime})>h(t)$, then by Condition $(2.c)$ of Definition 5.1, $L(t^{\prime})>L(t)$ except if the middle edge of all lozenges between $t$ and $t^{\prime}$ are colored $m$. In the latter case, let $s$ (resp. $s^{\prime}$) be the edge of type $\ell-1$ such that $t,s$ form a reverse triangle (resp. $t^{\prime},s^{\prime}$ form a direct triangle). Since $s_{\ell}=t_{\ell}+1$ and $s^{\prime}_{\ell}=t^{\prime}_{\ell}+1$, the inequality $t^{\prime}_{\ell}>t_{\ell}$ implies that $s^{\prime}_{\ell}>s_{\ell}$. Similarly, since $s_{\ell-1}=t_{\ell-1}-1$ and $s^{\prime}_{\ell-1}=t^{\prime}_{\ell-1}$, we have
@@
Hence, $L(s^{\prime})\leq L(s)$. Let us introduce the third edge $r$ (resp. $r^{\prime}$) of the triangle with edges $s,t$ (resp. $s^{\prime},t^{\prime})$. By Figure 12 and Condition (1) from Definition 5.1, we get that $e^{3}_{\ell-1}=N-e^{3}_{\ell}-e^{3}_{\ell+1}=N-L(t^{\prime})-L(r^{\prime})-1=L(s^{\prime})$ and $e^{2}_{\ell-1}=N-e^{2}_{\ell}-e^{2}_{\ell+1}=N-L(t)-L(r)=L(s)+1$ and thus $e^{2}_{\ell-1}=L(s)+1>L(s^{\prime})=e^{3}_{\ell-1}$, so that
@@
If $h(t)\geq h(t^{\prime})$ and $t_{\ell+1}>t^{\prime}_{\ell+1}$, then $t^{\prime}_{\ell+2}>t_{\ell+2}$. Suppose without loss of generality that $e^{1}_{\ell+1}\geq e^{2}_{\ell+1}$ and $e^{3}_{\ell+1}\geq e^{4}_{\ell+1}$. Let us consider the edges $s,s^{\prime}$ of type $\ell+1$ such that $(t,s,u)$ and $(t^{\prime},s^{\prime},u^{\prime})$ are respectively direct and reverse triangles of $T_{k}$ so that the corresponding edge of $t$ and the piece containing the direct triangle is $e^{2}$ and the corresponding edge for $t^{\prime}$ and the reverse triangle is $e^{3}$. Since $h(s)=t_{\ell+1}-1$ and $s_{\ell+2}=t_{\ell+2}+1$ and $h(s^{\prime})=t^{\prime}_{\ell+1}-1$ and $s^{\prime}_{\ell+2}=t^{\prime}_{\ell+2}$, $h(s)>h(s^{\prime})$ and $s^{\prime}_{\ell+2}\geq s_{\ell+2}$, so that $L(s)>L(s^{\prime})$ by Remark 5.2. Then, since $(t,u,s)$ is a direct triangle, Figure 12, $e^{2}_{\ell+1}=L(s)+1$, except if $c(t)=1,c(s)=0$ and $c(u)=m$ where $e^{2}_{\ell+1}=L(s)+2$. Likewise, since $(t^{\prime},s^{\prime},u^{\prime})$ is a reverse triangle, $e^{3}_{\ell+1}=L(s^{\prime})+1$ except if $c(t)=1,c(s)=0,c(u)=m$ or $c(t)=0,c(s)=m,c(u)=1$ where $e^{3}_{\ell+1}=L(s^{\prime})$. Hence, in any case,
@@
Lemma 5.11 (Different type blocks do not cross).
Let $t,t^{\prime}\in T_{k}$ be of respective type $\ell,\ell+1$ yielding edges in $\mathcal{E}$, and denote by $e^{1},e^{2}$ (resp. $e^{3},e^{4}$) the edges of $T_{N}$ corresponding to $t$ (resp. $t^{\prime}$). Then,
@@
The proof of the four assertions are similar.
@@
    Suppose that $t^{\prime}_{\ell}>t_{\ell}$ and $t^{\prime}_{\ell+1}<t_{\ell+1}$. Let $(t^{\prime},s^{\prime},r^{\prime})$ and $(t^{\prime},s^{\prime\prime},r^{\prime\prime})$ be the reverse and direct triangles belonging to pieces yielding respectively $e^{3}$ and $e^{4}$, with $s^{\prime},s^{\prime\prime}$ of type $\ell$. Since then $e^{4}_{\ell}\geq e^{3}_{\ell}$, it suffices to show that $e^{3}_{\ell}>e^{2}_{\ell}$. Since $s^{\prime}_{\ell}=t^{\prime}_{\ell}-1$ and $s^{\prime}_{\ell+1}=t^{\prime}_{\ell+1}+1$, we have $s^{\prime}_{\ell}\geq t_{\ell}$ and $t_{\ell+1}\geq s^{\prime}_{\ell+1}$, so that $L(s^{\prime})\geq L(t)$. Since $t^{\prime}$ is of type $\ell+1$ not and colored $m$ and $(t^{\prime},s^{\prime},r^{\prime})$ is a reverse triangle, Figure 12 and Condition (1) from Definition 5.1 yield that either $e^{3}_{\ell+2}=L(r^{\prime})+1$ and $L(r^{\prime})+L(s^{\prime})+L(t^{\prime})=N-2$ or $e^{3}_{\ell+2}=L(r^{\prime})$ and $L(r^{\prime})+L(s^{\prime})+L(t^{\prime})=N-1$. In any case, $e^{3}_{\ell+2}=N-1-L(t^{\prime})-L(s^{\prime})$, so that
    $e^{3}_{\ell}=N-e^{3}_{\ell+1}-e^{3}_{\ell+2}=N-L(t^{\prime})-(N-1-L(t^{\prime})-L(s^{\prime}))=L(s^{\prime})+1>L(t)=e^{2}_{\ell}=e^{1}_{\ell}.$
@@
    Suppose that $t^{\prime}_{\ell}<t_{\ell}$ and $t^{\prime}_{\ell+1}\geq t_{\ell+1}$, and let $s^{\prime}$ be the edge of type $\ell$ such that $(t^{\prime},s^{\prime},r^{\prime})$ is a direct triangle. Since $s^{\prime}_{\ell}=t^{\prime}_{\ell}$ and $s^{\prime}_{\ell+1}=t^{\prime}_{\ell+1}+1$, $s^{\prime}_{\ell}<t_{\ell}$ and $s^{\prime}_{\ell+1}>t_{\ell+1}$, so that $L(t)>L(s^{\prime})$ by Remark 5.2. Since $(t^{\prime},s^{\prime},r^{\prime})$ is a direct triangle, Figure 12 and Condition (1) from Definition 5.1 yield by a same reasoning as above that $e^{4}_{\ell+2}\geq L(r^{\prime})+1=N-L(s^{\prime})-L(t^{\prime})$, so that, using that $e^{4}_{\ell+1}=L(t^{\prime})$,
@@
    Suppose that $t^{\prime}_{\ell}\geq t_{\ell}$ and $t^{\prime}_{\ell+1}\geq t_{\ell+1}$, and let $s$ be the edge of type $\ell+1$ such that $(t,s,r)$ is a reverse triangle. Then, $s_{\ell}=t_{\ell}+1$ and $s_{\ell+1}=t_{\ell+1}-1$. Hence, $t^{\prime}_{\ell+1}>s_{\ell+1}$ and $t^{\prime}_{\ell+2}=N-t^{\prime}_{\ell+1}-t^{\prime}_{\ell}\leq N-s_{\ell+1}-1-s_{\ell}+1\leq s_{\ell+2}$ and the inequality is strict except when $t^{\prime}_{\ell}=t_{\ell}$ and $t^{\prime}_{\ell+1}=s_{\ell+1}$. Hence, by Remark 5.2 in the case of strict inequality and Condition (2.b) and (2.c) from Definition 5.1, $L(t^{\prime})>L(s)$, except when $t^{\prime}_{\ell}=t_{\ell},t^{\prime}_{\ell+1}=s_{\ell+1}$ and $C(r)=m$, in which case $L(t^{\prime})=L(s)$. In the first case, by Figure 12 we have $e^{2}_{\ell+1}\leq L(s)+1\leq L(t^{\prime})$. In the second case, since $C(r)=m$ we have $e^{2}_{\ell+1}=L(s)\leq L(t^{\prime})$, so that in any case